### v2

# arXiv:2504.09110v2

Abstract
Since many physical laws—from classical mechanics to electromagnetism—are formulated as two-body interactions, the same perspective naturally extends to biological and social dynamics. Here we focus on rhythmic phenomena, where phase reduction theory shows that synchronization dynamics can be universally described by coupled phase oscillators. Estimating the interaction functions of such systems from data offers a direct path to understanding and predicting such dynamics. Existing Fourier-series-based methods encounter difficulties with limited or biased data. To overcome this, we employ Gaussian process regression with additive periodic kernels. In our approach, we incorporate information about the estimation target into the statistical model in advance by designing kernel functions that capture the characteristics—additivity and $2\pi$-periodicity—of the coupling functions. Furthermore, owing to the Bayesian framework, our method enables the evaluation of uncertainty in the estimation results. We validate our approach on Van der Pol, FitzHugh-Nagumo, and spiking neural models. Our approach outperforms Fourier-series baselines in both error and stability under biased phase sampling. This enables data-driven studies of rhythm dynamics across a broader range of datasets. Furthermore, it makes a first step toward a statistically grounded, data-driven approach to general many-body systems with two-body interactions.
[…]
A55
1 Introduction
Figure 1: Conceptual overview. (Left) Not only basic physical laws such as classical mechanics, but also interactions in biological networks and rhythm synchronization can be formulated as the sum of two-body interactions. Rhythmic synchronization phenomena are universally described by coupled phase oscillators with a phase difference $\Delta\phi_{ij}$. (Right) The two-body coupling function is estimated by Gaussian process regression using an additive periodic kernel. Reflecting the phase bias reconstructed from the rhythmic time series, the uncertainty of the estimation results is expressed as the variance of the posterior distribution.
Interactions in many-body systems are often expressed as superpositions of two-body interactions (Fig. 1, Left). Typical examples include the gravitational force between point masses in classical mechanics and the Coulomb force between point charges in electromagnetism. Even when higher-order interactions exist in the system, it is meaningful to first approximate the system by considering only two-body interactions. This leads us to believe that, when analyzing unknown many-body systems such as communication among numerous organisms or active matter, it is reasonable to expect that the interactions between elements can be expressed as the sum of two-body interactions. If such interactions can be estimated from observational data, this provides a powerful means of uncovering underlying physical laws.
Numerous methods have been proposed for estimating interactions from observational data Galan2005 ; Tokuda2007 ; Stankovski2012 ; ota2014 ; Brunton2016 ; Raissi2018 ; Chen2018 ; Lu2021 . With recent advances in machine learning, it has become feasible to estimate even highly nonlinear interactions when sufficient data are available. In particular, kernel methods are effective for modeling nonlinear relationships and, when combined with Bayesian inference, enable quantification of the uncertainty in the estimated functions. However, to apply such approaches to the estimation of two-body interaction functions, the kernel must be expressible as a sum of two-body kernels. Generic kernels typically do not satisfy this decomposition, which limits their applicability in this context.
In this paper, we focus on rhythmic interactions, which are known to be reduced to systems of coupled phase oscillators Kuramoto1984 ; Winfree1980 ; Nakao2016 . This model has recently gained renewed prominence, as highlighted by the 2025 Boltzmann Medal awarded to Professor Yoshiki Kuramoto for his foundational contributions to the statistical physics of synchronization. Owing to its mathematical simplicity and universality, this model is widely employed to describe the rhythmic behavior of diverse complex systems, such as the coordinated activity of neurons in the brain Viriyopase2018 or the synchronized flashing of fireflies Ermentrout1991 . However, existing methods for estimating rhythm interactions ota2014 based on Fourier series face difficulties in accurately estimating the interaction function of oscillators in a synchronized state, due to the bias in the data. To overcome this difficulty, we employ Gaussian process regression, a nonparametric kernel method Rasmussen2006 , and propose a new approach that embeds the essential features of coupling functions in phase-oscillator models—namely periodicity and additivity—into the kernel design, introducing an additive periodic kernel (Fig. 1, Right).
This paper is organized as follows. In Section 2, we briefly review coupled phase-oscillator models. In Section 3, we propose a method for extracting the coupling functions of oscillator systems from data exhibiting rhythmic phenomena using Gaussian process regression. The proposed method is then applied to various systems in Section 4. In particular, we demonstrate the superiority of our approach over existing methods using coupled FitzHugh–Nagumo oscillators under conditions of biased or limited data. Furthermore, we apply the method to a many-body system of spiking neurons and confirm that all coupling functions are estimated with high accuracy. Finally, we conclude the paper and discuss further applications in Section 5.
2 Preliminaries
2.1 Coupled phase oscillators
Let us consider a system comprising $N$ interacting dynamical systems:
$\dot{\bm{X}_{i}}=F_{i}(\bm{X}_{i})+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}G_{j\to i}(\bm{X}_{i},\bm{X}_{j})\quad i=1,2,\cdots,N.$ (1)
Here, the function $F_{i}$ represents the spontaneous dynamics of the element $i$, and $G_{j\to i}$ represents a weak action from the element $i$ to the element $j$. We assume that each element has a stable limit cycle (i.e., each element is an oscillator). It is well known that such a system can be reduced to coupled phase oscillators Kuramoto1984 ; Winfree1980 .
$\dot{\phi_{i}}=\omega_{i}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\Gamma_{ij}(\Delta\phi_{ij}),\quad\Delta\phi_{ij}:=\phi_{j}-\phi_{i}.$ (2)
The functions $\{\Gamma_{ij}\}$ are called coupling functions, and for each $i$ and $j$, $\Gamma_{ij}$ represents the action from the oscillator $j$ to the oscillator $i$.
The advantage of employing the phase-oscillator model lies in the significant simplification of the system. While the original system (1) is generally multidimensional, the phase oscillators (2) assign only one variable $\phi_{i}$ to each limit cycle. Consequently, the focus can be directed solely towards rhythm synchronization characteristics.
2.2 estimation for coupling functions
In this paper, we estimate all phase coupling functions $\{\Gamma_{ij}\}$ from sampled phase time series $\{\bm{\phi}(\ell\delta t)\}_{\ell=1}^{n_{\mathrm{data}}}$ with $\bm{\phi}(\ell\delta t):=(\phi_{1}(\ell\delta t),\cdots,\phi_{N}(\ell\delta t))\in\mathbb{T}^{N}$ for $\ell=1,2,\ldots,n_{\mathrm{data}}$. Here, $\delta t$ denotes the sampling interval and $\mathbb{T}:=[0,2\pi)$. Since the phase coupling functions determine synchronization characteristics of the oscillators, it is desirable to estimate these functions from observed time series.
However, obtaining time-series data that comprehensively spans a high-dimensional state space is exceedingly challenging. Even if such data were acquired, the amount of data would be enormous, necessitating significant computational resources for estimation. To avoid this difficulty, we focus on the fact that the phase coupling functions are represented as the sum of two-body interactions, which enables us to use an additive kernel. See section 3.1 for details.
3 Methodology: Gaussian process regression
We employ the Gaussian process regression to estimate the coupling functions. In this section, we briefly address the procedure of the Gaussian process regression.
Suppose we have a time-series data $\mathcal{D}=\{(\bm{\phi}(\ell\delta t),\bm{y}(\ell\delta t))\}_{\ell=1}^{n_{\mathrm{data}}}$, which
$\bm{y}(\ell\delta t)$ $:=\left(\frac{\phi_{1}(\ell\delta t+\delta t)-\phi_{1}(\ell\delta t)}{\delta t},\cdots,\frac{\phi_{N}(\ell\delta t+\delta t)-\phi_{N}(\ell\delta t)}{\delta t}\right)\in\mathbb{R}^{N},$ (3)
and a function $\Gamma_{ij}\colon[0,2\pi)\to\mathbb{R}$, such that
$\begin{split}y_{i}&=f_{i}(\Delta\bm{\phi}_{i})+\xi_{i},\quad f_{i}(\Delta\bm{\phi}_{i})=\omega_{i}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\Gamma_{ij}(\Delta\phi_{ij}),\\ \Delta\bm{\phi}_{i}&:=(\Delta\phi_{i1},\cdots,\Delta\phi_{iN})\in\mathbb{T}^{N-1}\end{split}$ (4)
for $\ell=1,\dots,n_{\mathrm{data}}$, where $\xi_{i}$ is a random variable of $\mathcal{N}(0,\sigma^{2})$ that represents a noise in the output. Our task is to estimate the unknown functions $\{\Gamma_{ij}\}$ from the data $\mathcal{D}$ using the Gaussian process. If we estimate the sums $\{f_{i}\}$ of the two-body interactions, we can restore $\{\Gamma_{ij}\}$ except for the uncertainty of constants, so we will explain the estimation of $\{f_{i}\}$ below.
We start by assuming that the unknown regression function $f_{i}$ is drawn from a given Gaussian process prior,
$f_{i}\sim\mathcal{GP}(m_{i},k_{i}),$ (5)
where $m_{i}\colon\mathbb{T}^{N-1}\to\mathbb{R}$ is called a mean function and $k_{i}\colon\mathbb{T}^{N-1}\times\mathbb{T}^{N-1}\to\mathbb{R}$ is called a covariance function. The covariance function should be chosen so that it reflects one’s prior knowledge or belief about the regression function $f_{i}$, and we will discuss this in the next subsection. In many cases, the mean function $m_{i}$ is set to a constant zero function for simplicity.
The posterior distribution of $f_{i}$ is calculated analytically by linear algebra, and is again a Gaussian process. The distribution is given by the following closed form:
$f_{i}\mid\mathcal{D}\sim\mathcal{GP}(\overline{m}_{i},\overline{k}_{i}),$ (6)
where the posterior mean function $\overline{m}_{i}$ and the posterior covariance function $\overline{k}_{i}$ are
$\overline{m}_{i}(\Delta\bm{\psi})$ $=m_{i}(\bm{\psi})+k_{i;\bm{\psi}\Phi}(k_{i;\Phi\Phi}+\sigma^{2}I_{n_{\mathrm{data}}})^{-1}(\bm{y}_{i}-m_{i;\Phi}),$ (7)
$\overline{k}_{i}(\Delta\bm{\psi},\Delta\bm{\psi}^{\prime})$ $=k_{i}(\Delta\bm{\psi},\Delta\bm{\psi}^{\prime})-k_{i;\bm{\psi}\Phi}(k_{i;\Phi\Phi}+\sigma^{2}I_{n_{\mathrm{data}}})^{-1}k_{i;\Phi\bm{\psi}^{\prime}}.$ (8)
where
$[k_{i;\Phi\Phi}]_{mn}$ $=k_{i}(\Delta\bm{\phi}_{i}(m\delta t),\Delta\bm{\phi}_{i}(n\delta t)),$
$k_{i;\Phi\bm{\psi}}$ $=k_{i;\bm{\psi}\Phi}^{\top}=(k_{i}(\Delta\bm{\phi}_{i}(0),\Delta\bm{\psi}),\dots,k(\Delta\bm{\phi}_{i}(n_{\mathrm{data}}\delta t),\Delta\bm{\psi}))^{\top},$
$m_{i;\Phi}$ $=(m_{i}(\Delta\bm{\phi}_{i}(0)),\dots,m_{i}(\Delta\bm{\phi}_{i}(n_{\mathrm{data}}\delta t)))^{\top},$
$\bm{y}_{i}$ $=(y(\Delta\bm{\phi}_{i}(0)),\dots,y(n_{\mathrm{data}}\delta t))^{\top}.$
Gaussian process regression has the advantage that all calculations can be done in a closed form using only matrix operations. However, as the number of data $n_{\mathrm{data}}$ increases, inverse matrix calculations for $n_{\mathrm{data}}\times n_{\mathrm{data}}$ matrices are required, and the amount of memory and computation is enormous. Many sparse approximations have been proposed to overcome the computational complexity of Gaussian process regression fitc ; Quinonero2005 ; titsias09 ; hensman13 ; hensman15 .
3.1 Choice of the covariance function
The covariance function is a fundamental component of Gaussian process regression, as it encodes our assumptions about the function which we wish to predict (see kanagawa2018 for a detailed discussion). In our problem setting, since the coupling function have a domain $\mathbb{T}^{N-1}$, which means the function is $2\pi$-periodic, it is appropriate to use a periodic kernel as the covariance function. In addition, since the coupling function is additive (i.e., it is represented as the sum of two-body interactions), it can be seen that the kernel is also additive Duvenaud2011 . In conclusion, we propose to use following additive periodic kernel to estimate the coupling functions:
$k_{i}(\Delta\bm{\psi},\Delta\bm{\psi}^{\prime})=\sum_{j=1}^{N-1}\theta_{0}^{(j)}\exp\left(\theta_{1}^{(j)}\cos(\Delta\psi_{j}-\Delta\psi_{j}^{\prime})\right).$ (9)
Here, $\{\theta_{0}^{(j)}\}_{j=1}^{N-1}$ and $\{\theta_{1}^{(j)}\}_{j=1}^{N-1}$ are positive hyperparameters. $\theta_{0}^{(j)}$ determines the amplitude of the covariance function. $\theta_{1}^{(j)}$ can be said as an inverse lengthscale, which specifies the width of the $j$-axis direction of the covariance function.
3.2 Optimization
In the Gaussian process regression described so far, hyperparameters remain in the kernel $\theta_{0,1}$ and the noise strength $\sigma$. For brevity of notation, we will write these parameters as $\bm{\theta}$.
To estimate $\bm{\theta}$, we consider the maximum likelihood estimation, which makes inferences about the population that is most likely to have generated the data $\mathcal{D}$. The (marginal) log-likelihood for parameters $\bm{\theta}$ is
$\begin{split}\mathcal{L}(\bm{\theta})&:=\log p(\bm{y}\mid\bm{\phi},\bm{\theta})\\ &=-\frac{1}{2}\bm{y}^{\top}K_{\bm{\theta}}\bm{y}-\frac{1}{2}\log\det K_{\bm{\theta}}-\frac{n_{\mathrm{data}}}{2}\log 2\pi,\end{split}$ (10)
where $K_{\bm{\theta}}=k_{\Phi\Phi}+\sigma^{2}I_{n_{\mathrm{data}}}$ is the covariance matrix for the noisy data $\mathcal{D}$. The most probable parameters $\bm{\theta}$ are, therefore, estimated by finding the maximum of the log-likelihood function $\mathcal{L}(\bm{\theta})$.
The problem of finding the point that maximizes the marginal likelihood while changing hyperparameters is formulated as an optimization problem. In this paper, we use the gradient descent method is to find the maximum point of the marginal likelihood. The hyperparameters $\bm{\theta}$ is updated using the gradient with the following manner:
$\bm{\theta}^{(t+1)}\leftarrow\bm{\theta}^{(t)}+\alpha\frac{\partial\mathcal{L}}{\partial\bm{\theta}}(\bm{\theta}^{(t)}),$ (11)
where $\alpha$ denotes the learning rate.
The stochastic gradient descent (SGD) method, viewed as a stochastic approximation to standard gradient-based optimization, offers an alternative approach to optimization. In SGD, the gradient is estimated using a different minibatch of data at each iteration, and the parameters are updated accordingly. Its main advantage lies in reduced computational cost compared to full-batch gradient methods, which require evaluating the gradient using all data at every step. Moreover, the inherent stochasticity in minibatch selection helps the algorithm avoid getting trapped in local minima. Furthermore, the derivatives of the covariance function with respect to its hyperparameters can be computed analytically, enabling efficient optimization via automatic differentiation. Note that the implementation was performed using GPflow GPflow2017 .
4 Numerical Simulations
4.1 Coupled phase oscillators
Figure 2: Estimated all coupling functions (red) and true ones (black, dotted). Note that the estimated result is not strictly the coupling function itself, but the frequency added to it.
To confirm the validity of our approach, we first consider a simple 3-body system of coupled phase oscillators:
$\begin{aligned} &\dot{\phi}_{1}=1.0+0.01\sin(\phi_{2}-\phi_{1})-0.02\sin(\phi_{3}-\phi_{1}),\\ &\dot{\phi}_{2}=1.1+0.04\sin(\phi_{1}-\phi_{2})+0.01\sin(\phi_{3}-\phi_{2}),\\ &\dot{\phi}_{3}=0.95+0.01\sin(\phi_{1}-\phi_{3})-0.03\sin(\phi_{2}-\phi_{3}),\end{aligned}$ (12)
where $\phi_{1},\phi_{2},\phi_{3}$ are the phases of the oscillators. We prepare the time series by solving the ODEs (12) with forward Euler method from $t_{0}=0.0$ to $t_{1}=15.0$ using a randomly selected initial value and sample the data with sampling rate 0.2. We repeat the above calculation 50 times. Estimation results are shown in Fig. 2.
4.2 Coupled Van der Pol oscillators
Next, we apply the proposed method to coupled Van der Pol oscillators:
$\begin{split}&\dot{x}_{1}=y_{1}+K(x_{2}-x_{1})+\xi_{x_{1}}(t),\\ &\dot{y}_{1}=\varepsilon_{1}(1-x_{1}^{2})y_{1}-x_{1}+Kx_{2}^{2}y_{2}+\xi_{y_{1}}(t),\\ &\dot{x}_{2}=y_{2}-Kx_{1}^{2}y_{1}+\xi_{x_{2}}(t),\\ &\dot{y}_{2}=\varepsilon_{2}(1-x_{2}^{2})y_{2}-x_{2}+Kx_{1}y_{1}^{2}+\xi_{y_{2}}(t),\end{split}$ (13)
with $\langle\xi_{\alpha}(s)\xi_{\beta}(t)\rangle=\sigma^{2}\delta_{\alpha,\beta}\delta(s-t)$ for $\alpha,\beta\in\{x_{1},y_{1},x_{2},y_{2}\}$. The values of parameters are $\varepsilon_{1}=0.3,\varepsilon_{2}=0.7,K=0.01,\sigma=0.03$.
We reconstruct the phase time series from the obtained time-series data using the following method. First, we calculate the geometric angle $\vartheta_{i}(t)$ as follows:
$\vartheta_{i}(t)=\arg\left\{-\hat{x}_{i}(t)+\mathrm{i}\hat{y}_{i}(t)\right\},$ (14)
where $\hat{x}_{i}$ and $\hat{y}_{i}(t)$ are centered time series of $x_{i}(t)$ and $y_{i}(t)$, respectively. Even though $\vartheta_{i}(t)\in[0,2\pi)$ for all $t$, it is inappropriate to consider $\vartheta_{i}(t)$ as the phase time series of the oscillator $i$. This is because $\vartheta_{i}$ does not vary monotonically in the absence of interactions and noise. To overcome this difficulty, the following transformation from $\vartheta_{i}$ to $\theta_{i}$ should suffice Kralemann2007 ; Kralemann2008 :
$\theta_{i}(t)=2\pi\int_{0}^{\vartheta_{i}(t)}\sigma_{i}(\varphi)\mathrm{d}\varphi,$ (15)
where $\sigma_{i}$ denotes the empirical distribution of $\vartheta_{i}$. We consider $\theta_{i}(t)$ as the phase of oscillator $i$ and use them in the regression.
Estimation results are shown in Fig. 3. The results are highly consistent with the true coupling functions derived using the adjoint method. For comparison purposes, we also present the estimated coupling functions with the existing method ota2014 .
Figure 3: Estimation results for the coupling function in equation 13. Theory (dotted black curve), results obtained with existing method (solid blue curve) and with proposed method (solid red curve). The curve of the theory was obtained with the adjoint method. The highest order of the Fourier series was determined by the evidence approximation.
4.3 Coupled FitzHugh-Nagumo oscillators
Here is another example which demonstrates the effectiveness of the proposed method. Consider the following coupled FitzHugh-Nagumo oscillators FitzHugh1961 ; Nagumo1962 :
$\begin{split}&\dot{v}_{1}=v_{1}-\frac{1}{3}v_{1}^{3}-w_{1}+I_{\mathrm{ext}}+Kv_{1}v_{2}^{3}+\xi_{v_{1}}(t),\\ &\tau_{1}\dot{w}_{1}=v_{1}+a-bw_{1}+\xi_{w_{1}}(t),\\ &\dot{v}_{2}=v_{2}-\frac{1}{3}v_{2}^{3}-w_{2}+I_{\mathrm{ext}}+\xi_{v_{2}}(t),\\ &\tau_{2}\dot{w}_{2}=v_{2}+a-bw_{2}+\xi_{w_{2}}(t),\end{split}$ (16)
with $\langle\xi_{\alpha}(s)\xi_{\beta}(t)\rangle=\sigma^{2}\delta_{\alpha,\beta}\delta(s-t)$ for $\alpha,\beta\in\{v_{1},w_{1},v_{2},w_{2}\}$. Here, $a=1,b=0.8,I_{\mathrm{ext}}=0.8,K=0.0087,\tau_{1}=0.1^{-1},\tau_{2}=0.09^{-1}$.
We estimated the coupling functions for various numbers of data (Fig. 4). Specifically, we varied the number of initial values used to create the data (denoted by $N_{\text{init}}$) and the number of oscillator rounds per initial value used for estimation (denoted by $N_{\text{cycle}}$). We set the sampling rate as 0.2. The phase time series are reconstructed using the same method as described in section 4.2.
From Fig. 4, it is evident that when the data sample size is small, our proposed method accurately aligns with the theoretical result in estimating the coupling function, whereas the existing method fails in estimation. This implies that our method is superior to the other methods when a number of the data are small and the data are biased.
Figure 4: Estimated coupling function $\Gamma_{12}$ of the coupled FitzHugh-Nagumo system (16). Theory (dotted black curve), results obtained with existing method (solid blue line) and with proposed method (solid red line). (a) $N_{\text{init}}=10,N_{\text{cycle}}=5$. (b) $N_{\text{init}}=30,N_{\text{cycle}}=5$. (c) $N_{\text{init}}=100,N_{\text{cycle}}=5$. (d) $N_{\text{init}}=100,N_{\text{cycle}}=5$ when the part of the data where the phase difference is $0.5\pi$ or more and $\pi$ or less is discarded.
4.4 Spiking Neural Network oscillators
Finally, we consider a network of multiple coupled Hodgkin-Huxley models h&h and estimate the coupling function for all of them. There are seven neurons, five of which are excitatory and two of which are inhibitory. Details of the Hodgkin-Huxley model is summarized in Appendix A.
We employed a kernel function in the form of equation (9) with $N=7$ and $N=20$. See the lower graphs of Fig. 5 and Fig. 6 for the estimation result compared to the theoretical coupling functions obtained by the phase reduction approach. We confirm that the coupling functions between each oscillators are qualitatively consistent with the results from Gaussian process regression and the theoretical results.
Figure 5: (a) A network of spiking neurons coupled with five excitatory and two inhibitory neurons, (b) time series of the action potential of each neuron, (c) theoretically derived coupling functions between each neuron (black dotted lines) and coupling functions estimated from the data using Gaussian process regression (red solid lines: excitatory connection; blue solid lines: inhibitory connection; gray solid lines: no connection). Note that each coupling function $\Gamma_{ij}$ is translated to take 0 at $\Delta\phi_{ij}=0$ for comparison.
Figure 6: We apply the proposed method to a spiking neural network with 20 neurons (15 excitatory and 5 inhibitory). Theoretically derived coupling functions between each neuron (black dotted lines) and coupling functions estimated from the data using Gaussian process regression (red solid lines: excitatory connection; blue solid lines: inhibitory connection; gray solid lines: no connection). Note that each coupling function $\Gamma_{ij}$ is translated to take 0 at $\Delta\phi_{ij}=0$ for comparison.
5 Conclusions and Discussions
In this paper, we have proposed a new method for estimating the coupling functions of phase-oscillator models using Gaussian process regression. The kernel functions are designed to capture the characteristics of the coupling functions—additivity and periodicity. We have validated the effectiveness of the proposed method across various oscillator systems, including simple phase oscillators, the Van der Pol equations, and the FitzHugh–Nagumo equations. In particular, we have confirmed that stable estimation is achievable even when the data are biased or limited, with our method outperforming existing approaches. We also have demonstrated that the estimated coupling functions are highly consistent with theoretical ones in many-body system consisting of 20 spiking neurons.
The result in this paper opens new avenues for estimating interactions in many-body systems. Two points are particularly noteworthy. First, our method provides a practical guideline for designing kernel functions, which has long been a challenge in kernel methods. Second, by employing Gaussian process regression, a nonparametric Bayesian inference, our approach is not only robust against data bias but also enables the quantification of uncertainty in the estimation results.
In this study, we focused on rhythmic phenomena, i.e., systems with temporal periodicity. Nevertheless, our approach may also be applicable to systems with spatial periodicity. For example, in a school of swimming fish, as illustrated in the lower left of Fig. 1, interactions that align the orientation (angle) of the fish are expected to occur. In such cases, the interaction function is generally unknown, and researchers have typically assumed simple interactions, such as those in the Vicsek model Vicsek1995 , when constructing models. By contrast, our method enables the estimation of complex interactions directly from data, allowing for the development of models that more accurately reflect reality. In this sense, we believe our findings are highly valuable for applications to many-body systems with unknown interactions, such as active matter.
Acknowledgement
This work was supported in part by the following: MEXT KAKENHI Grant Numbers JP23H04467; JSPS KAKENHI Grant Numbers JP24H00723, JP22KJ1902, JP20K20520; JST BOOST Grant Number JPMJBS2407.
Appendix A Spiking neurons
We write below for details of the Hodgkin-Huxley model h&h .
$\begin{split}C\dot{V}=&G_{\mathrm{Na}}m^{3}h(E_{\mathrm{Na}}-V)+G_{\mathrm{K}}n^{4}(E_{K}-V)\\ &+G_{\mathrm{L}}(E_{\mathrm{L}}-V)+I_{\mathrm{input}}+\xi_{V},\\ \dot{m}=&\alpha_{m}(V)(1-m)-\beta_{m}(V)m+\xi_{m},\\ \dot{h}=&\alpha_{h}(V)(1-h)-\beta_{h}(V)m+\xi_{h},\\ \dot{n}=&\alpha_{n}(V)(1-n)-\beta_{n}(V)n+\xi_{n},\end{split}$ (17)
with the parameter values $C=1,G_{\mathrm{Na}}=120,G_{\mathrm{K}}=36,G_{\mathrm{L}}=0.3,E_{\mathrm{Na}}=50,E_{\mathrm{K}}=-77,E_{\mathrm{L}}=-54.4$. The auxiliary functions $\alpha_{m,h,n},\beta_{m,h,n}$ are
$\alpha_{m}(V)$ $=\frac{0.1(V+40)}{1-\exp[(-V-40)/10]},$
$\alpha_{h}(V)$ $=0.07\exp\frac{-V-65}{20},$
$\alpha_{n}(V)$ $=\frac{0.01(V+55)}{1-\exp[(-V-55)/10]},$
$\beta_{m}(V)$ $=4\exp\frac{-V-65}{18},$
$\beta_{h}(V)$ $=\frac{1}{1+\exp[(-V-35)/10]},$
$\beta_{n}(V)$ $=0.125\exp\frac{-V-65}{80},$
model of fast-spiking neurons,
$\begin{split}C\dot{V}=&G_{\mathrm{Na}}m^{3}h(E_{\mathrm{Na}}-V)+G_{\mathrm{K}}n^{2}(E_{K}-V)\\ &+G_{\mathrm{L}}(E_{\mathrm{L}}-V)+I_{\mathrm{input}}+\xi_{V},\\ \dot{m}=&\alpha_{m}(V)(1-m)-\beta_{m}(V)m+\xi_{m},\\ \dot{h}=&\alpha_{h}(V)(1-h)-\beta_{h}(V)m+\xi_{h},\\ \dot{n}=&\alpha_{n}(V)(1-n)-\beta_{n}(V)n+\xi_{n},\end{split}$ (18)
with the parameter values $C=1,G_{\mathrm{Na}}=112,G_{\mathrm{K}}=224,G_{\mathrm{L}}=0.1,E_{\mathrm{Na}}=55,E_{\mathrm{K}}=-97,E_{\mathrm{L}}=-70.0$. The auxiliary functions $\alpha_{m,h,n},\beta_{m,h,n}$ are
$\alpha_{m}(V)=$ $\frac{40(V-75)}{1-\exp[(75-V)/13.5]},$
$\alpha_{h}(V)=$ $0.0035\exp\frac{-V}{24.186},$
$\alpha_{n}(V)=$ $\frac{V-95}{1-\exp[(95-V)/11.8]},$
$\beta_{m}(V)=$ $1.2262\exp\frac{-V}{42.248},$
$\beta_{h}(V)=$ $\frac{0.017(-51.25-V)}{\exp[(-51.25-V)/5.2]-1},$
$\beta_{n}(V)=$ $0.025\exp\frac{-V}{22.222}.$
For each cell $i$, the input current is the sum of the bias and synaptic currents,
$I_{\mathrm{input},i}=I_{\mathrm{bias},i}+\sum_{j\in\mathrm{pre}_{i}}I_{\mathrm{syn},ij}.$ (19)
Here, $\mathrm{pre}_{i}$ denotes the set of indices of the cells that send synaptic inputs to the $i$th cell. We set the bias currents as $I_{\mathrm{bias},i}=30,32,6,6.5,34,36,38$ for $i=1,2,\dots,7$, respectively.
The current flowing through the synapses, $I_{\mathrm{syn},ij}$, is modeled using the kinetic synapse model Destexhe1994 , where it is represented as
$I_{\mathrm{syn},ij}=G_{ij}r_{ij}(t)[V_{i}(t)-E_{ij}].$ (20)
The fraction of bound receptor proteins is represented by $r_{i}j$, and its dynamics are described by the following equation:
$\frac{\mathrm{d}r_{ij}}{\mathrm{d}t}=\alpha_{ij}T_{ij}(1-r_{ij})-\beta_{ij}r_{ij},$ (21)
where $T_{ij}$ is the concentration of neurotransmitters, which is set to 1 when a spike is emitted by the presynaptic cell and resets to 0 after 1 millisecond. The constants $\alpha_{ij}$ and $\beta_{ij}$ govern the kinetics of $r_{ij}$, while $E_{ij}$ is the reversal potential and $G_{ij}$ is the synaptic conductance. The values used for excitatory and inhibitory synapses are $(\alpha_{ij},\beta_{ij},E_{ij},G_{ij})=(1.1,0.67,0,0.5)$ and (9.8,0.2,75,0.4)(9.8,0.2,-75,0.4) respectively. Additionally, a weak, independent noise function $\xi_{x,i}$ is added to the membrane voltage $V_{i}$ and channel variables $m_{i}$, $h_{i}$ and $n_{i}$. The noise follows a Gaussian white noise distribution, with $\braket{\xi_{x,i}(t)}=0$ and $\braket{\xi_{x,i}(t)\xi_{y,j}(s)}=\sigma^{2}_{x}\delta_{xy}\delta_{ij}\delta(t-s)$, where $x,y=V,m,h,n$, and $i$ and $j$ are the cell indices. The noise strengths used are $\sigma_{V}=0.5$ and $\sigma_{m}=\sigma_{h}=\sigma_{n}=5\times 10^{-6}$.